# The Structure of Hypergraphs Arising in Cellular Mobile Communication Systems

### v3

@@ 10 Extensions to Non-integral Values of the Path Loss Exponent $\alpha$
In typical wireless scenarios, the path loss exponent is between $1.6$ and $6.5$ [23]. In the present section, we prove some results that hold for all $\alpha\geq 1$.
@@
For all values of the path loss exponent $\alpha\geq 1$, the hypergraph $K_{1,6}$ is not realizable.
Proof: Fix $\alpha\geq 1$. By way of contradiction, suppose the hypergraph $H=K_{1,6}$ is generated by some wireless network $\langle S,\alpha,\beta\rangle$. By Lemma 12, it can be assumed without loss of generality that $\beta=1$. Let $S=\{s_{0},s_{1},\ldots,s_{6}\}$ be a collection of points in $\mathbb{R}^{2}$ such that the edge set $\mathcal{E}(H)$ is exactly $\{\{s_{0},s_{i}\}:i=1,\ldots,6\}$. In particular, $W=\{s_{1},\ldots,s_{6}\}$ is an independent set.
Translation in the Euclidean plane preserves distances, so it can be assumed without loss of generality that $s_{0}$ is at the origin. For $s_{i}\in W$, $\{s_{0},s_{i}\}$ is a forbidden set, whence $\frac{1}{d(s_{0},s_{i})}\geq 1$, or equivalently, $d(s_{0},s_{i})\leq 1$. Thus, the points in $W$ lie inside the (closed) unit ball. Because $W$ is an independent set, the distance between any two points in $W$ is greater than $1$. By Lemma 11, the points in $W$ can be pushed radially outward to the unit circle, and $W$ would remain an independent set.
@@
Let $r(\alpha)$ be a function from the domain $[1,\infty)$ to the positive integers, defined as follows: $r(\alpha):=$ the largest integer $k$ such that the hypergraph $K_{1,k}$ is generated by some wireless network with path loss exponent $\alpha$.
In other words, $r(\alpha)$ is the maximal value of $k$ such that the star hypergraph $K_{1,k}$ is realizable for the given value of $\alpha$. In [21], it was shown that when $\alpha=4$, the hypergraph $K_{1,4}$ is realizable and the hypergraph $K_{1,5}$ is not realizable. This implies that $r(\alpha)=4$. The results so far in the present paper imply that $r(1)=2$, $r(2)=3$, and $r(3)=4$, giving Table 2.
$\alpha$ 1 2 3 4
$r(\alpha)$ 2 3 4 4
Table 2: Values of $r(\alpha)$ known so far.
Observe that as $\alpha$ increases, $r(\alpha)$ stays the same or increases. Intuitively, we expect this behavior because, for larger values of the path loss exponent $\alpha$, the signal strength decreases more quickly with distance, and this reduces interference and can potentially allow for more stations to transmit simultaneously. In other words, we expect intuitively that $r(\alpha)$ will be a nondecreasing function of $\alpha$. This is proved formally next.
@@
The function $r(\alpha)$ is a nondecreasing function of $\alpha$.
Proof: Fix $\alpha\geq 1$. Suppose the hypergraph $K_{1,k}$ is generated by some wireless network having path loss exponent $\alpha$. It suffices to show that for each $\alpha^{\prime}>\alpha$, there exists a wireless network with path loss exponent $\alpha^{\prime}$ that generates $K_{1,k}$.
Fix $\alpha^{\prime}$ to be any value greater than $\alpha$. By hypothesis, there exists a wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ generating $H=K_{1,k}$. By Lemma 12, it can be assumed without loss of generality that $\beta=1$. Let $S=\{s_{0},s_{1},\ldots,s_{k}\}\subseteq\mathbb{R}^{2}$, where $s_{0}$ is the center vertex of the star $K_{1,k}$, the edge set of $H$ is $\{\{s_{0},s_{i}\}:i\in[k]\}$, and $W=\{s_{1},\ldots,s_{k}\}$ is an independent set for $\mathcal{A}$. Without loss of generality, it can be assumed that $s_{0}$ is located at the origin. Because $\{s_{0},s_{i}\}$ is a forbidden set $(i\in[k])$, $\frac{1}{d(s_{0},s_{i})^{\alpha}}\geq 1$, whence $d(s_{0},s_{i})\leq 1$. Thus, the points in $W$ lie inside the unit ball. The distance between any two points in $W$ is greater than $1$ since $W$ is an independent set. By Lemma 11, the points in $W$ can be pushed radially outward to the boundary of the unit ball (i.e. the unit circle), and $W$ would continue to remain feasible.
Thus, there exists a set $W$ of $k$ points on the unit circle that form an independent set when the path loss exponent is $\alpha$. In particular, for each $w\in W$, the energy at $w$ due to transmissions by the remaining stations in $W$ is less than $1$. That is,
$\sum_{s\in W-\{w\}}\frac{1}{d(s,w)^{\alpha}}<1,\forall w\in W.$
The distance between any two distinct points in $W$ is greater than $1$. For all $w\in W$ and for all $s\in W-\{w\}$, because $d(s,w)>1$, it follows that $d(s,w)^{\alpha^{\prime}}>d(s,w)^{\alpha}$. Thus,
$\sum_{s\in W-\{w\}}\frac{1}{d(s,w)^{\alpha^{\prime}}}<\sum_{s\in W-\{w\}}\frac{1}{d(s,w)^{\alpha}}<1,\forall w\in W.$
Hence, $W$ is an independent set for the wireless network $\langle S,\alpha^{\prime},1\rangle$. The points in $W$ lie inside the unit ball, whence $\{s_{0},s_{i}\}$ is a forbidden set for the wireless network $\langle S,\alpha^{\prime},1\rangle$. It follows that the hypergraph generated by the wireless network $\langle S,\alpha^{\prime},1\rangle$ is the $2$-uniform hypergraph $K_{1,k}$.
@@
For all values of the path loss exponent $\alpha\geq 5$, the hypergraph $K_{1,5}$ is realizable.
Proof: By Proposition 24 and Proposition 26, in order to prove the realizability of $K_{1,5}$ for all $\alpha\geq 5$, it suffices to prove the realizability of $K_{1,5}$ for $\alpha=5$. So let $\alpha=5$.
Consider the wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$, where $S=\{s_{0},s_{1},\ldots,s_{5}\}$, $\alpha=5$, $\beta=1$, $s_{0}$ is at the origin, and $W=\{s_{1},\ldots,s_{5}\}$ is a set of $5$ points placed uniformly on the unit circle in counterclockwise order. Let $s_{1}$ have polar coordinates $(1,0)$, and more generally, let $s_{i}$ have polar coordinates $(1,(i-1)\frac{2\pi}{5})$, for $i\in[5]$. Since $d(s_{0},s_{i})=1$, the subsets $\{s_{0},s_{i}\}$ are forbidden sets $(i\in[5])$. To show that the hypergraph generated by the wireless network $\mathcal{A}$ is $K_{1,5}$, it suffices to show that $W$ is an independent set, i.e. that
@@
Let $W=\{s_{1},\ldots,s_{5}\}$ be a set of $5$ stations placed uniformly on the unit circle. Thus, the polar coordinates of station $s_{i}$ are $(1,(i-1)\frac{2\pi}{5})$, for $i\in[5]$. Suppose the path loss exponent $\alpha=5$. Then, the energy $E(W-\{s_{1}\},s_{1})$ at $s_{1}$ due to transmissions by the remaining stations satisfies the inequality
@@
We have that $r(\alpha)=5$, for all $\alpha\geq 5$.
So far, analytical methods were used to obtain the value of $r(\alpha)$ for certain values of $\alpha$. We now conduct computer simulations to extend these results. To this end, a special case - where the stations are placed uniformly on the unit circle - will first be investigated. As we shall see, the values obtained from these simulations provide a tight lower bound for and usually the exact value of $r(\alpha)$.
Let $\mathcal{U}_{k}$ denote a set $\{s_{1},s_{2},\ldots,s_{k}\}$ of $k$ stations $(k\geq 1)$ placed uniformly on the unit circle, where the polar form of the location of $s_{i}$ is $s_{i}=\left(1,(i-1)\frac{2\pi}{k}\right)$, for $i\in[k]$. Recall that $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\alpha$ iff
$E(\mathcal{U}_{k}-\{s_{1}\},s_{1})=\sum_{i=2}^{k}\frac{1}{d(s_{i},s_{1})^{\alpha}}<1,$
@@
Let $u(\alpha)$ be a function from the domain $[1,\infty)$ to the positive integers, defined as follows: $u(\alpha)$ is defined to be the largest integer $k$ such that the set $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\alpha$. Thus, $u(5)=5$ by Lemma 28 and Proposition 24.
@@
The function $u(\alpha)$ is a nondecreasing function of $\alpha$.
Proof: Fix $\alpha,\alpha^{\prime},$ with $\alpha<\alpha^{\prime}$. Suppose $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\alpha$ and with reception threshold $\beta=1$. Then, for path loss exponent $\alpha^{\prime}$, we have
$E(\mathcal{U}_{k}-\{s_{1}\},s_{1})=\sum_{i=2}^{k}\frac{1}{d(s_{i},s_{1})^{\alpha^{\prime}}}.$
The Euclidean distance between any two distinct points in $\mathcal{U}_{k}$ is greater than $1$. For all $s_{i}\in\mathcal{U}_{k}-\{s_{1}\}$, because $d(s_{i},s_{1})>1$, it follows that $d(s_{i},s_{1})^{\alpha^{\prime}}>d(s_{i},s_{1})^{\alpha}$. Thus,
$\sum_{s_{i}\in\mathcal{U}_{k}-\{s_{1}\}}\frac{1}{d(s_{i},s_{1})^{\alpha^{\prime}}}<\sum_{s_{i}\in\mathcal{U}_{k}-\{s_{1}\}}\frac{1}{d(s_{i},s_{1})^{\alpha}}<1.$
It follows that $\mathcal{U}_{k}$ is an independent set for the wireless network $\langle\mathcal{U}_{k},\alpha^{\prime},1\rangle$.
Because $u(\alpha)$ is a nondecreasing function of $\alpha$, the bisection search algorithm can be used to compute the smallest value of $\alpha$ such that $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\alpha$, as shown in Algorithm 1. In this algorithm, the main function smallestAlphaUkIsFeasible takes as arguments a positive integer $k$ and an initial search space $[low,high]$ for bisection search, taken to be $[1,10]$ in our instances. This function makes calls to the boolean function isUkFeasible(alpha,k), which returns True iff $\sum_{i=2}^{k}\frac{1}{d(s_{i},s_{1})^{\alpha}}<1$, where the $s_{i}$’s are points spaced uniformly on the unit circle. When the main function is called with parameters $(3,1,10)$, $(4,1,10)$, and $(5,1,10)$, the values returned are 1.2618565559387207, 2.543102741241455, and 4.855621814727783, respectively. In the rest of this exposition, for simplicity we shall round these numerical values to three decimal places.
@@
Output: The smallest $\alpha$ such that $\mathcal{U}_{k}$ is feasible for the wireless network $\langle\mathcal{U}_{k},\alpha,1\rangle$
@@
[…]
@@
9
Algorithm 1 Bisection search to determine smallest $\alpha$ such that $\mathcal{U}_{k}$ is feasible
The function $u(\alpha)$ is a nondecreasing, integer-valued function and increases like the unit step function. Computer simulations using the bisection method give that $u(\alpha)$ increases from $2$ to $3$ when $\alpha$ is approximately $1.262$, increases from $3$ to $4$ when $\alpha$ is approximately $2.543$, and increases from $4$ to $5$ when $\alpha$ is approximately $4.856$. To summarize, we have that
$u(\alpha)=\begin{cases}2,~{}~{}~{}\mbox{if }\alpha\in[1,1.262)\\ 3,~{}~{}~{}\mbox{if }\alpha\in(1.262,2.543)\\ 4,~{}~{}~{}\mbox{if }\alpha\in(2.543,4.856)\\ 5,~{}~{}~{}\mbox{if }\alpha>4.856\end{cases}$
The connection between $u(\alpha)$ and the function of interest, namely $r(\alpha)$, is made next.
@@
We have that $u(\alpha)\leq r(\alpha)$, for all $\alpha\in[1,\infty)$.
Proof: Fix $\alpha\geq 1$ and let $k$ be any positive integer. Suppose $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\alpha$. By Lemma 12, without loss of generality it can be assumed that this wireless network has reception threshold $\beta=1$. It suffices to show that $r(\alpha)\geq k$. Let $S=\mathcal{U}_{k}\cup\{s_{0}\}$, where $s_{0}$ is at the origin. The hypergraph generated by the wireless network $\langle S,\alpha,1\rangle$ is the $2$-uniform hypergraph $K_{1,k}$. It follows that $r(\alpha)\geq k$.
The value of $r(\alpha)$ can now be obtained for various nonintegral values of $\alpha$. For $\alpha$ in the interval, $(1.262,2]$, $u(\alpha)=3$; by Lemma 31, $r(\alpha)\geq 3$. Furthermore, $r(2)=3$. Since $r(\alpha)$ is a nondecreasing function, $r(\alpha)\leq 3$, for all $\alpha\in(1.262,2]$. Together, these results imply that $r(\alpha)=3$ for all $\alpha\in(1.262,2]$. Similarly, one can put together the results on $u(\alpha)$ obtained above using computer simulations to determine $r(\alpha)$ for various nonintegral values of $\alpha$. We then obtain:
@@
$r(\alpha)=\begin{cases}2,~{}~{}~{}\mbox{if }\alpha=1\\ 3,~{}~{}~{}\mbox{if }\alpha\in(1.262,2]\\ 4,~{}~{}~{}\mbox{if }\alpha\in(2.543,4]\\ 5,~{}~{}~{}\mbox{if }\alpha>4.858\end{cases}$
We leave for further research the problems of determining $r(\alpha)$ in the intervals $(1,1.262)$, $(2,2.543)$ and $(4,4.858)$, and the problems of obtaining the transition values $1.262$, $2.543$, and $4.858$ exactly using analytical methods.

### v4

@@ 10 Extensions to Non-integral Values of the Path Loss Exponent $\gamma$
In typical wireless scenarios, the path loss exponent is between $1.6$ and $6.5$ [25]. In the present section, we prove some results that hold for all $\gamma>0$.
@@
For all values of the path loss exponent $\gamma>0$, the hypergraph $K_{1,6}$ is not realizable.
Proof: Fix $\gamma>0$. By way of contradiction, suppose the hypergraph $H=K_{1,6}$ is generated by some wireless network $\langle S,\gamma,\beta\rangle$. By Lemma 14, it can be assumed without loss of generality that $\beta=1$. Let $S=\{s_{0},s_{1},\ldots,s_{6}\}$ be a collection of points in $\mathbb{R}^{2}$ such that the edge set $\mathcal{E}(H)$ is exactly $\{\{s_{0},s_{i}\}:i=1,\ldots,6\}$. In particular, $W=\{s_{1},\ldots,s_{6}\}$ is an independent set.
Translation in the Euclidean plane preserves distances, so it can be assumed without loss of generality that $s_{0}$ is at the origin. For $s_{i}\in W$, $\{s_{0},s_{i}\}$ is a forbidden set, whence $\frac{1}{d(s_{0},s_{i})^{\gamma}}\geq 1$, or equivalently, $d(s_{0},s_{i})\leq 1$. Thus, the points in $W$ lie inside the (closed) unit ball. Because $W$ is an independent set, the distance between any two points in $W$ is greater than $1$. By Lemma 13, the points in $W$ can be pushed radially outward to the unit circle, and $W$ would remain an independent set.
@@
Let $r(\gamma)$ be a function from the domain $(0,\infty)$ to the positive integers, defined as follows: $r(\gamma):=$ the largest integer $k$ such that the hypergraph $K_{1,k}$ is generated by some wireless network with path loss exponent $\gamma$.
In other words, $r(\gamma)$ is the maximal value of $k$ such that the star hypergraph $K_{1,k}$ is realizable for the given value of $\gamma$. In Corollary 18, it was shown that when $\gamma=4$, the hypergraph $K_{1,4}$ is realizable and the hypergraph $K_{1,5}$ is not realizable. This implies that $r(4)=4$. The results presented in this paper also indicate that $r(3)=4$, $r(2)=3$, and $r(\gamma)=2$ for $\gamma\in(0,1]$. Additionally, Proposition 25 establishes that $r(\gamma)\leq 5$ for $\gamma>0$. These findings are summarized in Table 3.
$\gamma$ $(0,1]$ $2$ $3$ $4$ $[5,\infty)$
$r(\gamma)$ $2$ $3$ $4$ $4$ $\leq 5$
Table 3: Values of $r(\gamma)$ known so far.
Observe that as $\gamma$ increases, $r(\gamma)$ remains constant or increases. Intuitively, this behavior is expected because, with a higher path loss exponent $\gamma$, the signal strength diminishes more rapidly with distance, thereby reducing interference. This reduction in interference can potentially enable more stations to transmit simultaneously. Thus, we intuitively expect $r(\gamma)$ to be a nondecreasing function of $\gamma$. This is formally proved next. We begin by proving a stronger statement, assuming a reception threshold of $\beta=1$, which is sufficient to establish the desired result.
Lemma 27.
Let $\gamma^{\prime}>\gamma>0$. If the wireless network $\langle S,\gamma,1\rangle$ generates the hypergraph $K_{1,k}$, then the wireless network $\langle S,\gamma^{\prime},1\rangle$ generates the hypergraph $K_{1,k}$ for all $\gamma^{\prime}>\gamma$.
Proof: Suppose the wireless network $\langle S,\gamma,1\rangle$ generates the hypergraph $K_{1,k}$. In this hypergraph, adjacent nodes must be within a distance of at most 1, while nonadjacent nodes are separated by a distance greater than 1. Increasing the path loss exponent $\gamma$, which raises distance values to a higher power, only amplifies this distance-based relationship. Hence, the assertion follows.
@@
The function $r(\gamma)$ is a nondecreasing function of $\gamma$.
Proof: Fix $\gamma>0$. Suppose the hypergraph $K_{1,k}$ is generated by some wireless network having path loss exponent $\gamma$. It suffices to show that for each $\gamma^{\prime}>\gamma$, there exists a wireless network with path loss exponent $\gamma^{\prime}$ that generates $K_{1,k}$.
Fix $\gamma^{\prime}$ to be any value greater than $\gamma$. By hypothesis, there exists a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ generating $H=K_{1,k}$. By Lemma 14, it can be assumed without loss of generality that $\beta=1$. By Lemma 27, the hypergraph generated by the wireless network $\langle S,\gamma^{\prime},1\rangle$ is the hypergraph $K_{1,k}$.
@@
For all values of the path loss exponent $\gamma\geq 5$, the hypergraph $K_{1,5}$ is realizable.
Proof: By Proposition 25 and Proposition 28, in order to prove the realizability of $K_{1,5}$ for all $\gamma\geq 5$, it suffices to prove the realizability of $K_{1,5}$ for $\gamma=5$. So let $\gamma=5$.
Consider the wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$, where $S=\{s_{0},s_{1},\ldots,s_{5}\}$, $\gamma=5$, $\beta=1$, $s_{0}$ is at the origin, and $W=\{s_{1},\ldots,s_{5}\}$ is a set of $5$ points placed uniformly on the unit circle in counterclockwise order. Let $s_{1}$ have polar coordinates $(1,0)$, and more generally, let $s_{i}$ have polar coordinates $(1,(i-1)\frac{2\pi}{5})$, for $i\in[5]$. Since $d(s_{0},s_{i})=1$, the subsets $\{s_{0},s_{i}\}$ are forbidden sets $(i\in[5])$. To show that the hypergraph generated by the wireless network $\mathcal{A}$ is $K_{1,5}$, it suffices to show that $W$ is an independent set, i.e. that
@@
Let $W=\{s_{1},\ldots,s_{5}\}$ be a set of $5$ stations placed uniformly on the unit circle. Thus, the polar coordinates of station $s_{i}$ are $(1,(i-1)\frac{2\pi}{5})$, for $i\in[5]$. Suppose the path loss exponent $\gamma=5$. Then, the energy $E(W-\{s_{1}\},s_{1})$ at $s_{1}$ due to transmissions by the remaining stations satisfies the inequality
@@
We have that $r(\gamma)=5$, for all $\gamma\geq 5$.
So far, analytical methods were used to obtain the value of $r(\gamma)$ for certain values of $\gamma$. We now conduct computer simulations to extend these results. To this end, a special case - where the stations are placed uniformly on the unit circle - will first be investigated. As we shall see, the values obtained from these simulations provide a tight lower bound for and usually the exact value of $r(\gamma)$.
Let $\mathcal{U}_{k}$ denote a set $\{s_{1},s_{2},\ldots,s_{k}\}$ of $k$ stations $(k\geq 1)$ placed uniformly on the unit circle, where the polar form of the location of $s_{i}$ is $s_{i}=\left(1,(i-1)\frac{2\pi}{k}\right)$, for $i\in[k]$. Recall that $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\gamma$ iff
$E(\mathcal{U}_{k}-\{s_{1}\},s_{1})=\sum_{i=2}^{k}\frac{1}{d(s_{i},s_{1})^{\gamma}}<1,$
@@
Let $u(\gamma)$ be a function from the domain $(0,\infty)$ to the positive integers, defined as follows: $u(\gamma)$ is defined to be the largest integer $k$ such that the set $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\gamma$. Thus, $u(5)=5$ by Lemma 30 and Proposition 25.
@@
The function $u(\gamma)$ is a nondecreasing function of $\gamma$.
Proof: Fix $\gamma,\gamma^{\prime},$ with $\gamma<\gamma^{\prime}$. Suppose $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\gamma$ and with reception threshold $\beta=1$. Then, for path loss exponent $\gamma^{\prime}$, we have
$E(\mathcal{U}_{k}-\{s_{1}\},s_{1})=\sum_{i=2}^{k}\frac{1}{d(s_{i},s_{1})^{\gamma^{\prime}}}.$
The Euclidean distance between any two distinct points in $\mathcal{U}_{k}$ is greater than $1$. For all $s_{i}\in\mathcal{U}_{k}-\{s_{1}\}$, because $d(s_{i},s_{1})>1$, it follows that $d(s_{i},s_{1})^{\gamma^{\prime}}>d(s_{i},s_{1})^{\gamma}$. Thus,
$\sum_{s_{i}\in\mathcal{U}_{k}-\{s_{1}\}}\frac{1}{d(s_{i},s_{1})^{\gamma^{\prime}}}<\sum_{s_{i}\in\mathcal{U}_{k}-\{s_{1}\}}\frac{1}{d(s_{i},s_{1})^{\gamma}}<1.$
It follows that $\mathcal{U}_{k}$ is an independent set for the wireless network $\langle\mathcal{U}_{k},\gamma^{\prime},1\rangle$.
Because $u(\gamma)$ is a nondecreasing function of $\gamma$, the bisection search algorithm can be used to compute the smallest value of $\gamma$ such that $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\gamma$, as shown in Algorithm 1. In this algorithm, the main function smallestGammaUkIsFeasible takes as arguments a positive integer $k$ and an initial search space $[low,high]$ for bisection search, taken to be $[1,10]$ in our instances. This function makes calls to the boolean function isUkFeasible(gamma,k), which returns True iff $\sum_{i=2}^{k}\frac{1}{d(s_{i},s_{1})^{\gamma}}<1$, where the $s_{i}$’s are points spaced uniformly on the unit circle. When the main function is called with parameters $(3,1,10)$, $(4,1,10)$, and $(5,1,10)$, the values returned are 1.2618565559387207, 2.543102741241455, and 4.855621814727783, respectively. In the rest of this exposition, for simplicity we shall round these numerical values to three decimal places. The code used to generate these simulation results is available on GitHub at [22], allowing for easy reproduction and extension of the results.
@@
Output: The smallest $\gamma$ such that $\mathcal{U}_{k}$ is feasible for the wireless network $\langle\mathcal{U}_{k},\gamma,1\rangle$
@@
2       mid = (low + high) / 2
3       tolerance = 0.00001
4       while $|\mbox{high}-\mbox{low}|>\mbox{tolerance}$ do
@@
14
Algorithm 1 Bisection search to determine smallest value of the path loss exponent $\gamma$ such that $\mathcal{U}_{k}$ is feasible
The function $u(\gamma)$ is a nondecreasing, integer-valued function and increases like the unit step function. Computer simulations using the bisection method give that $u(\gamma)$ increases from $2$ to $3$ when $\gamma$ is approximately $1.262$, increases from $3$ to $4$ when $\gamma$ is approximately $2.543$, and increases from $4$ to $5$ when $\gamma$ is approximately $4.856$. It is clear that $\mathcal{U}_{2}$ is feasible for all $\gamma>0$. To summarize, we have that
$u(\gamma)=\begin{cases}2,~{}~{}~{}\mbox{if }\gamma\in(0,1.262)\\ 3,~{}~{}~{}\mbox{if }\gamma\in(1.262,2.543)\\ 4,~{}~{}~{}\mbox{if }\gamma\in(2.543,4.856)\\ 5,~{}~{}~{}\mbox{if }\gamma>4.856\end{cases}$
The transition value of $1.262$ obtained above using computer simulations can also be derived using analytical methods:
@@
Consider the wireless network $\langle\mathcal{U}_{3},\gamma,1\rangle$, where $\mathcal{U}_{3}=\{s_{1},s_{2},s_{3}\}$ is a set of three stations placed uniformly on the unit circle, $\gamma$ denotes the path loss exponent, and the reception threshold is $1$. Then, $\mathcal{U}_{3}$ is feasible if and only if $\gamma>2\log_{3}2\approx 1.262$.
Proof: Because the three stations are placed uniformly on the unit circle, the distance between any two distinct stations is $\sqrt{3}$. Choose a station from $\mathcal{U}_{3}$, say $s_{1}$. The interference at $s_{1}$ due to transmissions by $s_{2}$ and $s_{3}$ is
$E(\{s_{2},s_{3}\},s_{1})=\frac{1}{(\sqrt{3})^{\gamma}}+\frac{1}{(\sqrt{3})^{\gamma}}=\frac{2}{(\sqrt{3})^{\gamma}},$
which is less than $1$ if and only if $\gamma>2\log_{3}2$.
The connection between $u(\gamma)$ and the function of interest, namely $r(\gamma)$, is made next.
Lemma 34.
We have that $u(\gamma)\leq r(\gamma)$, for all $\gamma>0$.
Proof: Fix $\gamma>0$ and let $k$ be any positive integer. Suppose $\mathcal{U}_{k}$ is feasible for some wireless network with path loss exponent $\gamma$. By Lemma 14, without loss of generality it can be assumed that this wireless network has reception threshold $\beta=1$. It suffices to show that $r(\gamma)\geq k$. Let $S=\mathcal{U}_{k}\cup\{s_{0}\}$, where $s_{0}$ is at the origin. The hypergraph generated by the wireless network $\langle S,\gamma,1\rangle$ is the $2$-uniform hypergraph $K_{1,k}$. It follows that $r(\gamma)\geq k$.
The value of $r(\gamma)$ can now be obtained for various nonintegral values of $\gamma$. For $\gamma$ in the interval, $(1.262,2]$, $u(\gamma)=3$; by Lemma 34, $r(\gamma)\geq 3$. Furthermore, $r(2)=3$. Since $r(\gamma)$ is a nondecreasing function, $r(\gamma)\leq 3$, for all $\gamma\in(1.262,2]$. Together, these results imply that $r(\gamma)=3$ for all $\gamma\in(1.262,2]$. Similarly, one can put together the results on $u(\gamma)$ obtained above using computer simulations to determine $r(\gamma)$ for various nonintegral values of $\gamma$. We then obtain:
@@
$r(\gamma)=\begin{cases}2,~{}~{}~{}\mbox{if }\gamma\in(0,1]\\ 3,~{}~{}~{}\mbox{if }\gamma\in(1.262,2]\\ 4,~{}~{}~{}\mbox{if }\gamma\in(2.543,4]\\ 5,~{}~{}~{}\mbox{if }\gamma>4.858\end{cases}$
We leave for further research the problems of determining $r(\gamma)$ in the intervals $(1,1.262)$, $(2,2.543)$ and $(4,4.858)$, the question of whether equality holds in Lemma 34, and the problems of obtaining the transition values $2.543$ and $4.858$ exactly using analytical methods.